OpenStudy (anonymous):

Assume that for some function f(x), the expression f(x+h)-f(x) simplifies to -h/(x(x+h). Knowing this find f'(x). Then, what was f(x)? I know this has to do with the limit of a derivative but I'm unsure how to proceed.

myininaya (myininaya):

well have you found the limit as h->0 yet?

myininaya (myininaya):

remember \[\lim_{h \rightarrow 0}\frac{1}{h}(f(x+h)-f(x))=f'(x)\]

myininaya (myininaya):

you have \[f'(x)=\lim_{h \rightarrow 0}\frac{1}{h}(\frac{h}{x(x+h)})\]

myininaya (myininaya):

evaluate the limit